### v1

# Tracing the impacts of Mount Pinatubo eruption on global climate using spatially-varying changepoint detection

Abstract
Significant events such as volcanic eruptions can have global and long lasting impacts on climate. These global impacts, however, are not uniform across space and time. Understanding how the Mt. Pinatubo eruption affects global and regional climate is of great interest for predicting impact on climate due to similar events. We propose a Bayesian framework to simultaneously detect and estimate spatially-varying temporal changepoints for regional climate impacts. Our approach takes into account the diffusing nature of the changes caused by the volcanic eruption and leverages spatial correlation. We illustrate our method on simulated datasets and compare it with an existing changepoint detection method. Finally, we apply our method on monthly stratospheric aerosol optical depth and surface temperature data from 1985 to 1995 to detect and estimate changepoints following the 1991 Mt. Pinatubo eruption.
Keywords Aerosol optical depth; Bayesian hierarchical model; Spatially-varying Changepoint, Temperature, Volcanic eruption
1 Introduction
The Mount Pinatubo eruption in June 1991 is the largest volcanic eruption in recent history that injected nearly 20 megatons of sulfur dioxide into the atmosphere. The aerosol cloud resulting from the eruption encircled the globe for a few weeks, and global changes to the atmosphere, including a decrease in surface temperature in the northern hemisphere of about 0.5 C, were observed up to two years after the event (Self et al., 1996). Studying the impact of volcanic eruptions are of high interest to the climate community since they serve as natural examples of Stratospheric Aerosol Injection (SAI), a controversial proposed solar climate intervention to help cool the earth and reduce the impact of global warming through a global dimming effect (Robock, 2014). Consequently, we aim for a more comprehensive understanding of Pinatubo to shed new light on this contentious technology against climate change.
Scientists have studied how radiative forcing and climate were altered by the Mt. Pinatubo eruption through changes in aerosol optical depth. They found that the radiative forcing changes lasted for about 3 years (Stenchikov et al., 2009), leading to an observed heating in the stratosphere of 2-3K (Labitzke and McCormick, 1992) and cooling of global surface temperatures of about 0.4 K (Dutton and Christy, 1992; Thompson et al., 2009). Although there is a consensus that the eruption “caused" these global atmospheric changes, most published studies provided only qualitative evidence at a global scale. In this study, we attempt to quantitatively trace the impact of Pinatubo eruption on aerosol optical depth and temperatures at a finer scale.
Since climate impacts from a volcanic eruption, or a general SAI event, are expected to cause sudden changes in regional climate. It is natural to use changepoints to trace the impact of eruption at regional scale. Changepoint detection methods aim to identify significant shifts in the underlying distribution of the data and have been employed to understand the impact of significant events on environment. For instance, Hallema et al. (2017) used changepoint analysis to assess the impacts of wildland fires on watershed annual water yield. Additionally, Robbins et al. (2011) developed a new changepoint test method to identify changes in the tropical cyclone record in the North Atlantic Basin along with climate change over the period 1851–2008.
There have been many developments in changepoint detection in both univariate or multivariate settings, including two review articles (Reeves et al., 2007; Aminikhanghahi and Cook, 2017) and references therein. Additionally, methods such as pruned exact linear time (PELT) algorithm (Killick et al., 2012) and product partition model (PPM) (Barry and Hartigan, 1992) have gained considerable popularity in the changepoint literature. In situations where time series are indexed by spatial locations, the change behavior of time series that are geographically nearby is likely to be similar due to spatial correlation. For such data, the direct application of traditional changepoint detection methods to each time series might overlook this spatial feature and fail to exploit spatial correlation for improved changepoint detection and estimation.
To address this limitation, Majumdar et al. (2005) employed a Bayesian approach to detect various types of changepoints by incorporating spatial correlation. Additionally, Xuan and Murphy (2007) extended the PPM to account for the dependency structure in multivariate time series using sparse Gaussian graphical models. While these methods can handle spatial dependence among time series, they assume a common changepoint across all spatial locations, imposing a strong limitation on their application. Recently, a spatially varying changepoint model was proposed by Berchuck et al. (2019) to monitor glaucoma progression in visual field data, where multivariate conditional autoregressive priors are used to model the spatial dependencies in the changepoints and the mean parameters. However, their model assumes the data is temporally independent. Another related work is Wang et al. (2023), who developed a spatially varying changepoint estimation method for functional time series observed over a spatial domain.
The changes in regional climate caused by Pinatubo eruption are expected to exhibit similarity at nearby locations due to spatial dependency. However, these changes are anticipated to occur at different times due to the diffusion of the injected aerosols across the globe. In particular, we expect that the time of the change will increase as the spatial distance from the event location increases, again due to the diffusion of aerosols. These unique characteristics of our data render existing methods inadequate for our application. Consequently, we propose a novel space-time changepoint detection method that can identify when, where, and to what extent climate impacts have occurred following the 1991 eruption of Mt. Pinatubo while respecting the particular features of the SAI event. Our approach departs from conventional methods by modeling changepoints as a spatial process, where the expected value increases with the spatial distance from the event origin. By considering the spatial correlation of the data and the expected diffusion of the observed impact from the event location, our method is expected to be more effective in capturing the spatial patterns of climate impacts associated with the SAI events. Since our focus is soley on tracing the impact of the Mt. Pinatubo eruption, we examine at most one change after the eruption.
Our method extends the work of Majumdar et al. (2005) by allowing the changepoints to vary spatially. An additional advantage of our method is that both detection and estimation of changepoint are integrated in our approach. While most traditional methods focus on the mean shift, our method can detect changes in either the mean or variance and thus offers greater flexibility.
The rest of this paper is organized as follows. Section 2 describes the data and presents exploratory data analysis results. Section 3 introduces the model formulation and sampling procedure. In Section 4, we conduct extensive simulation studies to evaluate the performance of our proposed method. Section 5 demonstrates our methodology on global aerosol optical depth (AOD) and surface temperature data observed around the time of the Mt. Pinatubo eruption. Finally, in section 6 we provide a brief summary and discuss future potential work.
2 Data
As mentioned in the 1, notable changes in AOD and surface temperatures caused by the Pinatubo eruption has been documented. While aerosols rapidly spread around the globe and experienced a near immediate and steep increase in magnitude following the eruption, changes in surface temperature were more subtle and took years to manifest. We consider the AOD that measures the extinction optical thickness at 550nm, named “TOTEXTTAU” in Global Modeling and Assimilation Office (2015b) (GMAO), and the surface temperature, named “tavg1_2d_slv_Nx” in Global Modeling and Assimilation Office (2015a) (GMAO), to evaluate the impact of the eruption. We obtained both data sets from the Modern-Era Retrospective analysis for Research and Applications, Version 2 (MERRA-2) reanalysis.
The monthly AOD data ranges from January 1985 to December 1995, covering 132 time points in total. The data is over a $48\times 24$ grid and covers the entire globe. Following (Self et al., 1996), we exclude latitudes above $60^{\circ}$ and below $60^{\circ}$ in our analysis, resulting in $16\times 48=768$ spatial locations. Figure 1(a) shows the AOD data on these 768 spatial grids. In general, the aerosol data exhibits a clear jump after the eruption, followed by an approximately linear decay in time toward its previous level. However, not all locations precisely agree on the time of the jump, likely due to the aerosol diffusion process. To enhance our analysis, we first remove the seasonality by applying Seasonal-Trend decomposition using LOESS (STL) (Cleveland, 1990) to AOD data at each location. Then, we estimate the temporal trend as a linear function of time using the historical data up to June 1991, the month of the eruption. If the estimated trend parameter is statistically significant at a 0.05 level, we remove the estimated trend for both pre- and post-eruption data, detrending the time series before the changepoint to ensure a mean of zero. Finally, we normalize the time series at each location to achieve approximately constant variance across all locations.
Fig. 1: Time Series of (a) stratospheric AOD data and (b) surface temperature data before preprocessing. Each blue curve represents a time series at a grid point. The red curve represents the global average.
The monthly surface temperature data, shown in Figure 1(b), also spans from January 1985 to December 1995 and is on the same $48\times 24$ grid as the AOD data. In contrast to the AOD data, the impact of volcanic eruption on temperatures is indiscernible through simple visualization, suggesting a more subtle influence on temperatures, if any. When examining how temperature responds to a volcanic eruption, it is customary to study the temperature zonal mean indexed by latitude (Robock and Mao, 1995; Stenchikov et al., 2002; Gao et al., 2008). Hence, we average temperatures over longitude for each of the 24 latitude bands given in the data and conduct the analysis on the latitudinal means. This reduces the number of observations to $24$, with each latitude representing a latitude band covering $7.5^{\circ}$. We employed the same procedure as the AOD data to detrend and normalize temperature data. The processed zonal mean temperatures can be seen in Figure 8.
To evaluate our conjecture that changes in these climate series vary spatially, we conducted exploratory data analysis on the global AOD data. At each grid location, we applied the Bayesian information criteria-based chagepoint detection procedure, which is a well-established method for climate series (Reeves et al., 2007). Figure 2 illustrates the heatmap of detected changepoints and estimated mean shifts at each grid location. Overall, the changepoints appear to be spatially varying and clustered, as does the magnitude of change. Although a few locations show changepoints prior to 1991/06 that are apparently not due to Mt. Pinatubo eruption, in general, the changepoints appear to spread latitudinally from the source of the eruption. This spatial pattern is consistent with what we expect from an SAI event, given the influence of strong zonal winds in the lower stratosphere that circulate the globe (Robock and Matson, 1983). These results not only verify our conjecture but also provide guidance for the development of our methodology.
Fig. 2: (a) Heatmap of detected changepoints for AOD. Darker color indicates earlier change. (b) Heatmap of estimated change amount. Darker color indicates stronger magnitude of mean shift. The red diamond marks the location of Mt. Pinatubo.
3 Method
We propose a Bayesian hierarchical model to detect and estimate spatially varying changepoints for a spatiotemporal data. We model the likelihood of data in the first level, provide the prior for latent processes in the second level, and close the hierarchy by specifying priors and hyperpriors for all unknown parameters in the model. We then discuss the challenges of MCMC for the proposed model.
3.1 Changepoint detection and estimation
Let $Y(\mathbf{s},t)$ denote the observed climate variable at location $\mathbf{s}\in\mathbb{R}^{2}$ and time $t\in\mathbb{N}$, $N$ be the total number of spatial locations, and $M$ the total time points. We propose the following model for the spatiotemporal process $Y(\mathbf{s},t)$ with possible mean and variance change in time:
Level 1 The likelihood of the $Y(\mathbf{s},t)$ process:
$Y(\mathbf{s},t)$ =$=\begin{cases}\mu_{1}(\mathbf{s},t)+U(\mathbf{s},t)+\epsilon_{1}(\mathbf{s},t),&t\leq\tau(\mathbf{s})\\ \mu_{2}(\mathbf{s},t)+U(\mathbf{s},t)+\epsilon_{2}(\mathbf{s},t),&t>\tau(\mathbf{s}),\end{cases}$ (1)
where $\tau(\mathbf{s})$ is the changepoint at location ${\bf s}$, $\mu_{1}(\mathbf{s},t)$ and $\mu_{2}(\mathbf{s},t)$ are the mean functions before and after changepoint, $U(\mathbf{s},t)$ is a zero mean space-time correlated error process, and $\epsilon_{1}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}N(0,\sigma_{1}^{2})$ and $\epsilon_{2}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}N(0,\sigma_{2}^{2})$ are the before and after changepoint measurement errors. When $\tau(\mathbf{s})=M$, it indicates the absence of changepoint at ${\bf s}$ as the entire process is governed by a single model. To keep our model parsimonious and avoid potential identifiablity issue, we assume a common spatiotemporal process $U(s,t)$ for both before and after changepoint. This ensures that all changes will be attributed to mean and/or variance of measurement errors. This assumption also allows us to take advantage of the space-time separablility of the covariance structure to facilitate the computation for spatiotemporal data. While Majumdar et al. (2005) allows for a change in the covariance structure by modeling the pre-changepoint error process as $U({\bf s},t)+V({\bf s},t)$ and the post-changepoint error process as $U({\bf s},t)+W({\bf s},t)$, their model assumes $\tau({\bf s})$ as well as $\mu({\bf s},t)$ to be a constant as a sacrifice to reduce the identifiability risk. Since the primary interest for our data is to identify the spatially varying changepoints while respecting the spatio-temporally varying mean, we opt for a model that provides this flexibility.
As mentioned in Section 2, the data after preprocessing can be assumed to have constant mean prior to the changepoint and a linear trend in time after changepoint. We also observed from Figure 2 that the change amount not only varies spatially but also tends to be clustered in space. To account for these features, we model the pre and post-changepoint mean processes as
$\mu_{1}(\mathbf{s},t)$ $=\alpha_{0}$ (2)
$\mu_{2}(\mathbf{s},t)$ $=\alpha_{0}+\gamma_{0}(\mathbf{s})+\gamma_{1}(\mathbf{s})\cdot(t-\tau(\mathbf{s}))$ (3)
$=\alpha_{0}+(\gamma_{0F}+\gamma_{0R}(\mathbf{s}))+(\gamma_{1F}+\gamma_{1R}(\mathbf{s}))\cdot(t-\tau(\mathbf{s})).$ (4)
In (2) and (3), $\alpha_{0}$ is the constant parameter representing the average global mean before the changepoint, $\gamma_{0}({\bf s})$ is the mean shift at the time of the changepoint at location ${\bf s}$, and $\gamma_{1}({\bf s})$ represents the slope after changepoint at locations ${\bf s}$. We further decompose $\gamma_{0}({\bf s})$ and $\gamma_{1}({\bf s})$ into a fixed and random components in (4). The parameters $\gamma_{0F}$ and $\gamma_{1F}$ can be interpreted as the average mean shift and post-changepoint temporal trend across all locations that have a changepoint. The parameters $\gamma_{0R}({\bf s})$ and ${\gamma_{1R}({\bf s})}$ serve as spatial random effects to allow the mean shift and trend to vary spatially. To respect the spatial correlation of change amount and the post-changepoint trend, we model $\gamma_{0R}({\bf s})$ and ${\gamma_{1R}({\bf s})}$ as a spatially correlated process.
Let ${\bf s}_{0}$ denote the location at which the change is first observed, i.e., ${\bf s}_{0}=\arg\min_{\bf s}\tau({\bf s}).$ Assuming that the change diffuses from the event origin, ${\bf s}_{0}$ can also be interpreted as the event location, i.e., Mt. Pinatubo in our context. We model $\tau(\mathbf{s})$ as the sum of two terms: (1) the changepoint at the event location ${\bf s}_{0}$, denoted by $\tau_{0}$, and (2) the temporal lag between changepoints at $\mathbf{s_{0}}$ and ${\bf s}$, denoted by $\Delta(\mathbf{s})$. Specifically, we have
$\tau(\mathbf{s})=\min\{M,\tau_{0}+\Delta(\mathbf{s})\}.$ (5)
We require $\Delta(\mathbf{s})\geq 0$ such that $\tau_{0}$ is the earliest changepoint, which ensures that all other changepoints detected are at or after $\tau_{0}$ and thus not due to other unrelated events. We also fix $\Delta(\mathbf{s}_{0})=0$ in our estimation so that $\tau(\mathbf{s}_{0})=\tau_{0}$ is observed as the earliest changepoint, i.e., $\tau(\mathbf{s}_{0})=\min_{{\bf s}}\tau({\bf s})$. Note that $\tau_{0}$ takes discrete values in $\{1,\ldots,M\}$ and $\tau_{0}=M$ implies that there is no changepoint at any of the locations. There are two main advantages of separating $\tau_{0}$ from $\tau(\mathbf{s})$. First, it ensures that the changepoint at any location $\mathbf{s}\neq\mathbf{s}_{0}$ occurs strictly after the changepoint at $\mathbf{s}_{0}$, as any changepoint prior to $\tau_{0}$ would be due to another event that is not of our interest. Second, we can model $\Delta(\mathbf{s})$ as a continuous process with a spatial correlation structure to reflect the spatial pattern observed in Figure 2(a) while keeping $\tau_{0}$ as a discrete variable. This allows for a convenient way to test for the existence of the changepoint and also incorporate the prior knowledge about the possible range of date at which the change was first observed from the eruption.
We assume $\boldsymbol{\Delta}=(\Delta({\bf s}_{1}),\ldots,\Delta({\bf s}_{N}))^{T}$ follow a log normal process with mean $\mathbf{X}\boldsymbol{\beta}$ and covariance matrix ${\bf\Sigma}_{\Delta}$, where $\mathbf{X}$ is the $N\times 2$ matrix of spatial lags from each location $\mathbf{s}$ to the event location ${\bf s}_{0}$, with the $i$th row of $\boldsymbol{X}$ given by $(\hbox{lon}(\mathbf{s}_{i})-\hbox{lon}(\mathbf{s}_{0}),\hbox{lat}(\mathbf{s}_{i})-\hbox{lat}(\mathbf{s}_{0}))$. $\boldsymbol{\beta}$ is the 2-dimensional vector with each component representing the rate of diffusion in longitude and latitude, respectively.
In other words, we model $\Delta({\bf s})$ as a spatial process that increases with the distance from $s_{0}$ to reflect the nature of the physical process we observe, i.e. the time it takes for the impact of the event to reach a specific location will tend to increase as its distance from the event location increases.
Let $\boldsymbol{\gamma}_{0R}$ and $\boldsymbol{\gamma}_{1R}$ be the vector $\gamma_{0R}({\bf s}_{i})$ and ${\gamma_{1R}({\bf s}_{i})}$ for $i=1,\ldots,N$, respectively. Let $\mathbf{U}$ be the vector stacking up all $U({\bf s}_{i},t_{j})$ for $i=1,\ldots,N$ and $j=1,\ldots,M$, and let $\log(\boldsymbol{\Delta})$ be the vector of element-wise logarithm of $\boldsymbol{\Delta}$. In our next level of the hierarchical model, we provide priors for the latent processes in model (1).
Level 2 Latent processes:
$\boldsymbol{\gamma}_{0R}$ $\sim N(\mathbf{0},\mathbf{\Sigma}_{\gamma_{0}}),$
$\boldsymbol{\gamma}_{1R}$ $\sim N(\mathbf{0},\mathbf{\Sigma}_{\gamma_{1}})$
$\mathbf{U}$ $\sim N(\mathbf{0},\mathbf{\Sigma_{U}}),$
$\log(\boldsymbol{\Delta})$ $\sim N(\mathbf{X}\boldsymbol{\beta},\mathbf{\Sigma}_{\Delta}).$
For computational simplicity, we assume that $\boldsymbol{\Sigma_{U}}$ has a separable space-time covariance structure with exponential covariance functions in both space and time, i.e. $\boldsymbol{\Sigma_{U}}=\sigma^{2}_{U}\mathbf{R}(\phi_{U})\otimes\mathbf{R}(\psi_{U}),$ where the $(i,j)$th element of $\mathbf{R}(\theta)$ is $R(\theta)_{ij}=\exp(-\theta h)$ for a distance $h$ either in space or time, and $\otimes$ is the Kronecker product. Since our data spans the entire globe, we employed the great circle distance as the distance metric for the spatial correlation matrix. While a positive definite covariance function defined on Euclidean space may not necessarily be valid on spheres, the exponential covariance function has been shown to be valid on spheres when the Euclidean distance is replaced with the great circle distance. (Gneiting, 2013; Huang et al., 2011). We assume $\mathbf{\Sigma}_{\gamma_{0}}=\sigma^{2}_{\gamma_{0}}\mathbf{R}(\psi_{\gamma_{0}})$, $\mathbf{\Sigma}_{\gamma_{1}}=\sigma^{2}_{\gamma_{1}}\mathbf{R}(\psi_{\gamma_{1}})$ and $\boldsymbol{\Sigma}_{\Delta}=\sigma^{2}_{\Delta}\mathbf{R}(\psi_{\Delta})$, where $\mathbf{R}(\psi_{\gamma_{0}})$, $\mathbf{R}(\psi_{\gamma_{1}})$ and $\mathbf{R}(\psi_{\Delta})$ are also governed by an exponential covariance function with great circle distance.
3.2 MCMC Sampling
To describe the details of our sampling, we first write Model (1) in a matrix form. Let $\mathbf{1}_{\tau}^{+}$ be a $MN$-length vector of 0’s and 1’s, where 1 indicates that the corresponding time index in $Y({\bf s}_{i},t_{j})$ is greater than $\tau({\bf s}_{i}).$ Let $\mathbf{1}_{\tau}^{-}=\mathbf{1}_{MN}-\mathbf{1}_{\tau}^{+}$ be the vector of pre-changepoint indices. Then, Model (1) can be written as
$\mathbf{Y}$ $=\boldsymbol{\mu}_{\tau}+\mathbf{U}+\boldsymbol{\epsilon}_{1}\circ\mathbf{1}_{\tau}^{-}+\boldsymbol{\epsilon}_{2}\circ\mathbf{1}_{\tau}^{+},$ (6)
where
$\boldsymbol{\mu}_{\tau}$ $=\boldsymbol{\mu}_{1}\circ\mathbf{1}_{\tau}^{-}+\boldsymbol{\mu}_{2}\circ\mathbf{1}_{\tau}^{+},$ (7)
where $\circ$ denotes the Schur product operator, and $\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2},\boldsymbol{\epsilon}_{1},\boldsymbol{\epsilon}_{2}$ denote the $MN$-length vector of $\mu_{1}({\bf s}_{i},t_{j})$, $\mu_{2}({\bf s}_{i},t_{j})$, $\epsilon_{1}({\bf s}_{i},t_{j})$ and $\epsilon_{2}({\bf s}_{i},t_{j})$ for $i=1,\ldots,N$ and $j=1,\ldots,M$. Let $\boldsymbol{\tau}$ be the vector of $\begin{pmatrix}\tau({\bf s}_{1}),\ldots,\tau({\bf s}_{N})^{T}\end{pmatrix}$. The full conditional likelihood of $\bf Y$, $f(\mathbf{Y}\lvert\mathbf{U},\boldsymbol{\tau},\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2},\sigma_{1}^{2},\sigma_{2}^{2})$, is
$f(\mathbf{Y}|\cdot)\propto|{\bf\Sigma}_{Y}|^{-1/2}\exp\left(-\frac{1}{2\sigma_{1}^{2}\sigma_{2}^{2}}(\mathbf{Y}-\boldsymbol{\mu}_{\tau}-\mathbf{U})^{T}{\bf\Sigma}_{Y}^{-1}(\mathbf{Y}-\boldsymbol{\mu}_{\tau}-\mathbf{U})\right),$ (8)
where ${\bf\Sigma}_{Y}=\sigma_{2}^{2}\text{diag}(\mathbf{1}_{\tau}^{-})+\sigma_{1}^{2}\text{diag}(\mathbf{1}_{\tau}^{+})$ is a diagonal matrix of pre and post-changepoint measurement variances. Alternatively, integrating out $\mathbf{U}$ gives us
$f(\mathbf{Y}|\cdot)\propto|{\bf\Sigma}_{Y}|^{-1/2}\exp\left(-\frac{1}{2\sigma_{1}^{2}\sigma_{2}^{2}}(\mathbf{Y}-\boldsymbol{\mu}_{\tau})^{T}({\bf\Sigma}_{Y}+{\bf\Sigma}_{U})^{-1}(\mathbf{Y}-\boldsymbol{\mu}_{\tau})\right).$ (9)
There are two main challenges in MCMC sampling for our model. First, the Appendix shows that obtaining the posterior for $\mathbf{U}$ involves inverting the covariance matrix ${\bf\Sigma}_{Y}^{-1}+{\bf\Sigma}_{U}^{-1}$. Although $\bf\Sigma_{U}$ is separable, (8) shows that the matrix ${\bf\Sigma}_{Y}$ involves vectors $\mathbf{1}_{\tau}^{-}$ and $\mathbf{1}_{\tau}^{+}$, which depends on $\tau({\bf s})$ and thus cannot be separated into a Kronecker product of $M\times M$ and $N\times N$ matrices. This significantly amplifies the computational cost associated with sampling $\mathbf{U}$. Secondly, sampling both $\mathbf{U}$ and $\boldsymbol{\mu}_{\tau}$ can in practice lead to an identifiability issue, since both $\mathbf{U}$ and $\boldsymbol{\mu}_{\tau}$ have the potential to capture the trend in the data. While integrating out $\mathbf{U}$ from the model could solve these problems, retaining $\mathbf{U}$ is more desirable from a computational perspective. This stems from the fact that $\mathbf{Y}$ is no longer conditionally independent when the conditioning on $\mathbf{U}$ is removed, as indicated by the non-diagonal covariance matrix ${\bf\Sigma}_{Y}+{\bf\Sigma}_{U}$ in (9). Consequently, sampling other parameters whose full conditionals involve the likelihood of $\mathbf{Y}$ becomes significantly more intricate.
To address the first issue, we propose a simple but effective approach. We assume, without the loss of generality, that the variance increases after the changepoint and write the post-changepoint measurement error as $\boldsymbol{\epsilon}_{2}=\boldsymbol{\epsilon}_{1}+\boldsymbol{\epsilon}_{\gamma},$ where $\boldsymbol{\epsilon}_{\gamma}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}N(0,\sigma^{2}_{\gamma})$ is a vector of white noise independent of $\boldsymbol{\epsilon}_{1}.$ Then, we can rewrite $f(\boldsymbol{Y}\lvert\cdot)$ in (8) as
$f(\boldsymbol{Y}\lvert\cdot)\propto\sigma_{1}^{MN}\exp\left(-\frac{1}{2\sigma_{1}^{2}}(\mathbf{Y}-\boldsymbol{\mu_{\tau}}-\mathbf{U}-\boldsymbol{\epsilon_{\gamma}}\circ\mathbf{1_{\tau}^{+}})^{T}(\mathbf{Y}-\boldsymbol{\mu_{\tau}}-\mathbf{U}-\boldsymbol{\epsilon_{\gamma}}\circ\mathbf{1_{\tau}^{+}})\right)$ (10)
and add $\boldsymbol{\epsilon_{\gamma}}\sim N(\mathbf{0},\sigma^{2}_{\gamma})$ to the level 2 prior. The full conditional for $\mathbf{U}$ now involves the covariance matrix $(\sigma^{2}_{1}\mathbf{I}_{MN})^{-1}+\bf\Sigma_{U}^{-1},$ which can be factored into a Kronecker product of a spatial and a temporal covariance matrix. If the variance decreases after the changepoint, we simply let $\boldsymbol{\epsilon}_{1}=\boldsymbol{\epsilon}_{2}+\boldsymbol{\epsilon}_{\gamma}$ where $\boldsymbol{\epsilon}_{\gamma}$ now is independent with $\boldsymbol{\epsilon}_{2}$. Then, we replace $\boldsymbol{\epsilon_{\gamma}}\circ\mathbf{1_{\tau}^{+}}$ with $\boldsymbol{\epsilon_{\gamma}}\circ\mathbf{1_{\tau}^{-}}$ and $\sigma^{2}_{1}$ with $\sigma^{2}_{2}$ in (10). One caveat of this approach is that we need to pre-specify whether the variance increases or decreases beforehand. We recommend trying both scenarios in parallel and choosing a better model based on Deviance Information Criterion (DIC).
To address the identifiability between $\bf U$ and $\boldsymbol{\mu}_{\tau}$ in sampling, we adopt the Partially Collapsed Gibbs (PCG) sampler proposed by Van Dyk and Park (2008), which allows us to remove the conditioning on $\mathbf{U}$ when sampling $\boldsymbol{\tau}$ and $\boldsymbol{\alpha}=(\alpha_{0},\gamma_{0},\gamma_{1})^{T}$. We refer to Van Dyk and Park (2008) and Van Dyk and Jiao (2015) for more details on how to derive the PCG sampler from an ordinary Gibbs sampler. Figure 3 illustrates the sampling procedure for our model under the PCG sampler. It is important to note that, unlike ordinary Gibbs samplers, the order of draws in the PCG sampler must be maintained, as permuting the order may alter the stationary distribution of the chain in the PCG sampler (Van Dyk and Jiao, 2015).
Ordinary Gibbs Sampler Step 1: $\tau\sim p(\tau|Y,U^{\prime},\theta^{\prime},\alpha^{\prime})$ Step 2: $\theta\sim p(\theta|Y,U^{\prime},\alpha^{\prime},\tau)$ Step 3: $\alpha\sim p(\alpha|Y,U^{\prime},\theta,\tau)$ Step 4: $U\sim p(U|Y,\alpha,\tau,\theta)$ Step 1: $\tau\sim p(\tau|Y,\theta^{\prime},\alpha^{\prime})$ Step 2: $\alpha\sim p(\alpha|Y,\theta^{\prime},\tau)$ Step 3: $U\sim p(U|Y,\tau,\alpha,\theta^{\prime})$ Step 4: $\theta\sim p(\theta|Y,U,\tau,\alpha)$ Partially Collapsed Gibbs Sampler
Fig. 3: Sampling procedure for Partially Collapsed Gibbs (PCG) sampler. In our context, $p(\cdot)$ denotes the target posterior distribution and $\theta=(\sigma^{2}_{1},\sigma^{2}_{2},\sigma^{2}_{\Delta},\boldsymbol{\epsilon_{\gamma}},\psi_{U},\phi_{U},\psi_{\Delta},\psi_{\gamma_{0}},\psi_{\gamma_{1}},\boldsymbol{\beta})^{T}.$ Compared to the ordinary Gibbs sampler, the PCG sampler drops the conditioning on $\mathbf{U}$ when sampling $\tau$ and $\alpha$, and then modifies the order of sampling in the ordinary Gibbs to ensure the stationary distribution of the chain under PCG.
The following discusses the sampling for each parameter in turn. The full conditionals for $\mathbf{U}$ and $\boldsymbol{\epsilon_{\gamma}}$ are $MN$-dimensional multivariate normal distributions. To sample $\mathbf{U}$, we leverage the compatibility of a Kronecker product plus a constant diagonal matrix with the eigenvalue decomposition (Stegle et al., 2011), which allows us to sample $\mathbf{U}$ using eigendecompositions of $M\times M$ and $N\times N$ matrices instead of having to factorize a $MN\times MN$ covariance matrix. While the posterior for $\boldsymbol{\epsilon_{\gamma}}$ is no longer identically distributed, the Appendix shows that its independence structure is preserved. Thus, posterior samples can be drawn easily even when the spatial and/or temporal dimension is high. See Appendix for more details.
Parameters $\boldsymbol{\Delta},\boldsymbol{\gamma_{0R}},$ and $\boldsymbol{\gamma_{1R}}$ are sampled via a Metropolis Hastings algorithm. While the full conditionals for $\boldsymbol{\gamma_{0R}}$ and $\boldsymbol{\gamma_{1R}}$ have closed forms, the resulting distribution involves $MN\times MN$ covariance matrices, making the sampling computationally expensive. To achieve efficient mixing, we employ an adaptive component-wise Metropolis-within-Gibbs (Roberts and Rosenthal, 2009) to propose a new state for each of the $N$ locations separately and automatically tune the acceptance ratios for each location as closely as possible to 0.44, a ratio considered optimal for one-dimensional Gaussian proposals (Roberts and Rosenthal, 2001).
The parameters $\sigma^{2}_{U},\sigma_{1}^{2},\sigma_{2}^{2},\sigma^{2}_{\Delta},\sigma^{2}_{\gamma_{0}},\sigma^{2}_{\gamma_{1}},\boldsymbol{\beta},$ and $\boldsymbol{\alpha}=(\alpha_{0},\gamma_{0},\gamma_{1})^{T}$ can be assigned conjugate priors and estimated via a Gibbs sampler. We specify weakly independent normal prior for $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$ and inverse gamma priors for variance parameters $\sigma^{2}_{1},\sigma^{2}_{2},\sigma^{2}_{U},$ and $\sigma^{2}_{\Delta}.$ For $\sigma^{2}_{\gamma_{0}}$ and $\sigma^{2}_{\gamma_{1}},$ we use a truncated inverse gamma prior and choose the upper truncation points to be a small fraction of $\gamma_{0F}$ and $\gamma_{1F}.$ This encourages the magnitude of random effect to be smaller than that of the fixed effect, which further ensures identifiability. The parameters $\psi_{U},\phi_{U},\psi_{\Delta},\psi_{\gamma_{0}},\psi_{\gamma_{1}}$ are given uniform priors and are sampled via the Metropolis Hastings algorithm. For $\tau_{0},$ any discrete probability distribution that takes values in $\{1,2,\ldots,M\}$ can be used as a prior. The full conditional distributions and sampling details are outlined in the Appendix.
4 Simulation Study
We perform simulation studies to evaluate the accuracy of our model in detecting three different types of changes – mean shift only, variance shift only, and mean and variance shift. We also explore how the magnitude of change affects the model’s performance.
4.1 Setup
To generate data for simulation, we select $N=121$ spatial locations on an $11\times 1$ 11 grid and randomly choose one of the grid locations as $\mathbf{s_{0}}$. At each location, we consider a time series of length $M=61$ and generate $\tau(\mathbf{s})$ using (5), with $\tau_{0}=18,\boldsymbol{\beta}=(1.5,1)^{T},$ $\sigma^{2}_{\Delta}=1,$ and $\psi_{\Delta}=0.5$. These values were chosen such that the number of locations with changepoint |$|\{{\bf s}:\tau(\mathbf{s})<M\}|$ is comparable to the number of locations without changepoints |$|\{{\bf s}:\tau(\mathbf{s})\geq M\}|$ on average. Since a changepoint detection is equivalent to a classification problem, having a balance between the two classes ensures our evaluation focuses on method performance without being affected by class imbalance. Without loss of generality, we fix $\alpha_{0}=0$ and $\sigma^{2}_{1}=1$. We let $\gamma_{0}$ take values in $\{0,1.5,2,3\}$ and $\sigma^{2}_{2}$ in $\{3,5,10\}$ to study the model performance under different strengths of mean and variance shift. The parameters $\sigma^{2}_{U},$ $\phi_{U},$ and $\psi_{U}$ are set to 1, 1.5, and 2, respectively, to emulate spatial and temporal dependence observed in real data based on our exploratory data analysis of global surface temperature data. We use model (1) to generate data and run 100 simulations for each setting. The ratio of number of locations without changepoints to those with changepoints in our simulated data ranges from 32:89 to 101:20, with an average of 57:64.
In implementing our method, we assume no prior knowledge about $\tau_{0}$ and use a discrete uniform prior. This is slightly more challenging than our real problem, since for the Pinatubo data, we know $\tau_{0}$ is likely right after the eruption. We give weakly informative prior for other unknown parameters. We use the posterior mode of $\min(M,\tau({\bf s}))$ as the estimated changepoint for location ${\bf s}$. If the mode for $\min(M,\tau({\bf s}))$ is at $M$, we determine that there is no changepoint at ${\bf s}$.
To evaluate model performance, we use the metrics of false positive rate (FPR) and false negative rate (FNR) to measure the accuracy of the changepoint detection, and use root mean squared error (RMSE), empirical coverage of the 95% credible interval (CI), and length of the CI to assess the changepoint estimation. FPR is defined as the ratio of the number of falsely detected locations to the total number of locations that have changepoints detected, and FNR is the ratio of the number of falsely undetected locations to the total number of locations without changepoints detected. When calculating RMSE, we treat the true changepoint as having value M (i.e., $\tau(\mathbf{s})=M$) when there is no changepoint at location ${\bf s}$. A combination of a narrower credible interval with empirical coverage closer to the nominal level indicates a more precise uncertainty quantification.
We compare the FPR, FNR and RMSE of our spatial-temporal model with the univariate time series changepoint detection method offered by the R package changepoint. The latter method is applied to each of the spatial locations separately (hereinafter 1D method). We use cpt.mean, cpt.var, and cpt.meanvar functions for each type of change, all using modified Bayesian information criteria penalty for estimating at most one changepoint (Hinkley, 1970).
4.2 Results
Figure 4 summarizes the results for detecting and estimating changepoints only in the mean under different magnitudes of change. Mean shift of zero indicates no changepoint, i.e. $\tau(\mathbf{s})=M$ for all locations, while a larger mean shift indicates a stronger signal.
Our method consistently returns a much lower FPR than the 1D method in all cases. When a changepoint is present, the FPR of our method decreases as the signal increases, while the FPR of the 1D method stays almost constant regardless of the signal strength. FNR of the two methods is comparable when the signal is strong, but our method largely outperforms 1D under weaker signal. FNR is not defined when there is no changepoint. Both the 1D and our method achieve smaller RMSE when signal is stronger, but our method outperforms the 1D method across all signal strengths with significantly smaller RMSE. The advantage of our method in terms of RMSE is more pronounced when the signal is weaker. Since the 1D method does not provide confidence intervals, we are unable to compare the two methods in terms of their uncertainty quantification. However, the empirical coverage of our 95% credible interval seems to be close to the nominal level for all tested cases, though their median coverage seems slightly above the nominal level. Although there is no simple interpretation for the empirical coverage of Bayesian credible intervals as there is for confidence intervals, our results still suggest a positive view of our credible intervals. The lengths of the credible interval appear to be informative given that the maximum interval length is $60$. Unsurprisingly, the length of the credible interval gets wider as signal becomes weaker.
Figure 5 shows the results for both variance shift only and mean and variance shift combined. We set $\gamma_{0}$ to be 0 and 2 associated with zero and positive mean shift, and set $\sigma^{2}_{\gamma}$ to be 3 and 5, representing two different levels of variance shift. Again, our method significantly outperforms the 1D method in both the detection and estimation of changepoints. The 1D method appears to be very conservative in detecting variance only shift, resulting in a low FPR but an extremely high FNR. This implies that the 1D method is more sensitive to mean changes than variance changes. However, when testing for a combined mean and variance shift, the 1D method detects changepoints for all locations and tends to identify a changepoint at $t=M-1$ when there is actually no changepoint, resulting in a low FNR but an extremely high FPR. In contrast, our method consistently performs well in detecting changepoints both in variance only change and in mean and variance change. The RMSE, empirical coverage, and length of credible intervals show a similar pattern as in Figure 4. Once again, both the RMSE and the length of credible intervals decrease as the signal increases. The results for mean and variance shift seem to be very similar to the corresponding mean shift only results with the mean shift set to 2. This suggests that the changepoint detection and estimation by our model are driven mainly by the mean shift rather than the variance shift in the presence of both types of changes.
Overall, our method outperforms the 1D method by taking advantage of the spatial correlation in the changepoint process for spatially indexed time series. The importance of borrowing strength from neighboring locations becomes more pronounced and necessary, especially when the change signal is weaker.
Fig. 4: Boxplots of RMSE, FPR, FNR, the empirical coverage probability and the length of 95% credible intervals under different mean shift signal strength. “ST” is our Spatial-Temporal model and “1D” is the univariate method from R package changepoint.
Fig. 5: Boxplots of RMSE, FPR, FNR, the empirical coverage probability and the length of 95% credible intervals under four settings comprised of two $\gamma_{0}$ values representing mean change and two $\sigma^{2}_{\gamma}$ values representing variance change. $\gamma_{0}=0$ indicates no mean change. “ST” is our Spatial-Temporal model and “1D” is the univariate method from R package changepoint.
5 Impact of Mt Pinatubo Volcanic Eruption
We apply our proposed method to AOD and temperature data to detect changes that are possibly caused by the stratospheric aerosol injection from the Mt. Pinatubo eruption in June 1991.
5.1 Stratospheric aerosol data
Following the suggestions in Section 3.2, we applied Model (1) on the aerosol data with both $\sigma^{2}_{1}<\sigma^{2}_{2}$ and $\sigma^{2}_{1}>\sigma^{2}_{2}$ since we have no prior knowledge about whether the variance increases or decreases after changepoint. In addition, we also test the case where $\sigma^{2}_{1}=\sigma^{2}_{2},$ which means that the change is soley due to a mean shift. We set ${\bf s}_{0}$ to be at ($123.4375^{\circ}$, $19^{\circ}$), which is an average of the AOD on the grid surface bounded by $(120^{\circ}$, $127.5^{\circ}$) in longitude and $(15^{\circ}$, $22.5^{\circ}$) in latitude. Since this grid contains the location of Mt. Pinatubo at ($120^{\circ}$, $15^{\circ}$), we anticipate the change to approximately spread from this location under our model assumption. Therefore, ${\bf s}_{0}$ is expected to have the earliest changepoint among all grid points, satisfying the definition ${\bf s}_{0}=\arg\min_{{\bf s}}\tau({\bf s})$. We used a discrete uniform prior for $\tau_{0}$ as well as weakly informative priors for all other parameters. Table 1 compares the DIC value and the posterior mean of key parameter estimates for the three forms of variance shift models. We observe that the equal variance model achieves the lowest DIC value. Consistent with this finding, the estimated variance shift parameter $\sigma^{2}_{\gamma}$ for the variance increase/decrease assumptions is very small compared to $\sigma^{2}_{1}$/$\sigma^{2}_{2}$. Thus, we proceed our inference based on the assumption $\sigma_{1}^{2}=\sigma_{2}^{2}$, that is, the change is solely due to a mean shift.
$\sigma^{2}_{1}=\sigma^{2}_{2}$ $\sigma^{2}_{1}<\sigma^{2}_{2}$ $\sigma^{2}_{1}>\sigma^{2}_{2}$ 95 % CI
DIC 3,526,069 3,638,721 3,719,177 for $\sigma^{2}_{1}=\sigma^{2}_{2}$
$\alpha_{0}$ 0.2754 0.1913 0.2499 (0.0754,0.4808)
$\gamma_{0F}$ 4.2408 4.2825 4.923 (3.4503,4.7342)
$\gamma_{1F}$0.0820 -0.0888 -0.0881 (-0.0996,-0.0646)
$\sigma^{2}_{U}$ 1.1449 1.1391 1.1410 (1.0131,1.3127)
$\sigma^{2}_{1}$ 0.3512 0.2790 0.1409 (0.2563,0.5245)
$\sigma^{2}_{2}$ NA 0.3092 0.1348 NA
$\sigma^{2}_{\gamma}$ NA 0.0302 0.0061 NA
$\beta_{1}$0.0186 -0.0092 -0.0151 (-0.1488,0.1020)
$\beta_{2}$ 0.5404 0.4447 0.4490 (0.1562,0.9623)
Table 1: DIC values and posterior summary of key parameters under different model assumptions for AOD data. The credible intervals in the last column are derived under $\sigma^{2}_{1}=\sigma^{2}_{2}$.
We report the 95% credible intervals of key parameters under the assumption of $\sigma_{1}^{2}=\sigma_{2}^{2}$ in Table 1. The significantly positive estimate of $\gamma_{0F}$ indicates an elevated aerosol level immdediately after the volcanic eruption, and the negative estimate for $\gamma_{1F}$ indicates the subsequent restoration of aerosol level post-eruption, consistent with the anticipated trend. The diffusion parameter for longitude, $\beta_{1}$, is not significantly different from 0. In contrast, there is strong evidence to show that the diffusion parameter for latitude, $\beta_{2}$, is substantially greater than 0. This observation aligns with the spatial pattern of changepoints predominantly driven by latitude, as depicted in Figure 2(a) and 6.
We determine the estimated changepoints by taking the posterior mode of the floor function of $\min(M,\tau(\mathbf{s}))$ at each location. Figure 6(a) shows the heatmap of estimated changepoints. Our model detects a changepoint at all locations, with estimated values ranging from Jun 1991 to Sep 1991. The estimated changepoints show a pattern driven more by latitude than longitude, with the earliest changepoints predominantly occurring along latitudes $3.5^{\circ}$ through $34^{\circ}$. This pattern aligns with our exploratory data analysis and is consistent with existing literature, which reports that the Pinatubo aerosol layer circled the Earth in 21 days and had spread to latitudes around $30^{\circ}N$ and $10^{\circ}S$ in the same period (Self et al., 1996). McCormick and Veiga (1992); Stowe et al. (1992) also found the Pinatubo aerosols straddled the equator.
Figure 6(b) displays the estimated mean shift parameter, $\gamma_{0}({\bf s})$, which ranges from $+0.25$ to $+1.5$. The spatial pattern of the estimated mean change matches with that shown in Figure 2(b), indicating the successful capture of the spatial variability in mean change by our method.
The time series of aerosol after preprocessing and the estimated $\mu_{1}$ and $\mu_{2}(\mathbf{s},t)$ from our model are shown in Figure 7. Our mean estimates exhibit the same pattern as the preprocessed data, where the time series close to equator jump first and tend to have a larger peak compared to those around the latitudes further north or south. These results, together with Figure 6, demonstrate that our model effectively captures the diffusion of the impact of the volcanic eruption on the AOD data. Specifically, the estimated changepoints for locations near the latitude of Mt. Pinatubo ($15^{\circ}$) coincide with the month of the eruption, while the changepoints for locations further north or south occur several months later. We also plot the residuals for the AOD series by location in the Supplement, which verifies that $\sigma_{1}^{2}=\sigma_{2}^{2}$ is a reasonable assumption.
Fig. 6: (a) Heatmap of detected changepoints. Darker color indicates earlier change. (b) Heatmap of estimated change amount. Darker color indicates larger magnitude of change. The red diamond marks the location of Mt. Pinatubo.
Fig. 7: Time series of (a) aerosol data after preprocessing and (b) posterior mean of $\mu_{1}$ and $\mu_{2}(\mathbf{s},t)$, zoomed in to years $1991-1993$ The color represents the latitude of their location. The blue, yellow and red vertical lines mark the average value of the changepoints in latitude bands $(60^{\circ}S-20^{\circ}S),(20^{\circ}S-20^{\circ}N),$ and $(20^{\circ}N-60^{\circ}N)$.
5.2 Latitudinal surface temperature data
Unlike the clear signal of volcanic impact on stratospheric AOD data, the impact on surface temperature is visually indiscernible, as observed in Figure 1. Indeed, the impact was too subtle to be detected when we applied our method to the gridded temperature data as was done for AOD data. Due to the zonal feature of the stratospheric aerosols spreading, the impact of volcanic eruption on surface temperature is more noticeable in latitude bands, as mentioned in Section 2. Consequently, we use zonal mean temperatures to trace the impact of eruption.The mechanism by which aerosols ejected into the lower stratosphere affect surface temperature is a complex process. Unlike AOD, we do not expect the volcanic impact on surface temperatures to diffuse from the event location Robock and Matson (1983). Thus, we negate the influence of the distance from $\mathbf{s_{0}}$ on $\tau_{0}({\bf s})$ by setting $\boldsymbol{\beta}=0$ in the mean of $\log(\boldsymbol{\Delta})$.
Previous literature has found that the effect of the Mt. Pinatubo eruption on surface temperature mostly took place in the two-year period following the eruption (Robock and Matson, 1983; Self et al., 1996). Therefore, we give $\tau_{0}$ an equal weight discrete uniform prior covering the months in the following two years and the last month in the data: $\{1991/06,1991/07,\ldots 1993/06,1995/12\}$. The last time point, $1995/12$, indicates the possibility that there is no observed changepoint at any of the locations.
Similar to aerosol data, we apply Model (1) with $\sigma^{2}_{1}<\sigma^{2}_{2}$, $\sigma^{2}_{1}>\sigma^{2}_{2}$, and $\sigma^{2}_{1}=\sigma^{2}_{2}$ on surface temperature and compare the results of the three model in Table 2. Again, the mean shift only model ($\sigma^{2}_{1}=\sigma^{2}_{2}$) achieves the lowest DIC, and the estimated variance change parameter $\sigma^{2}_{\gamma}$ for the two variance shift models is very small compared to $\sigma^{2}_{1}$/$\sigma^{2}_{2}$. We thus proceed our analysis under the assumption $\sigma^{2}_{1}=\sigma^{2}_{2}$. Table 2 reports the posterior mean and the 95% credible interval of key parameters. Unlike the aerosol data, the mean shift parameter $\gamma_{0F}$ for temperature is significantly negative, indicating the global cooling effect of eruption consistent with other literature. The slope parameter $\gamma_{1F}$ is also significantly negative, indicating continued cooling of the global temperature during the time period considered in our data.
$\sigma^{2}_{1}=\sigma^{2}_{2}$ $\sigma^{2}_{1}<\sigma^{2}_{2}$ $\sigma^{2}_{1}>\sigma^{2}_{2}$ 95% CI
DIC -1,692 -1,543 -1,631 for $\sigma^{2}_{1}=\sigma^{2}_{2}$
$\alpha_{0}$ 0.2092 0.2117 0.2103 (0.1182,0.3074)
$\gamma_{0F}$0.8440 -0.8592 -0.8430 (-1.1714,-0.5410)
$\gamma_{1F}$0.0117 -0.0116 -0.0116 (-0.0213,-0.0028)
$\sigma^{2}_{U}$ 0.7918 0.7933 0.7901 (0.7242,0.8655)
$\sigma^{2}_{1}$ 0.1029 0.1046 0.1068 (0.0798,0.1436)
$\sigma^{2}_{2}$ NA 0.1127 0.1011 NA
$\sigma^{2}_{\gamma}$ NA 0.0081 0.0057 NA
Table 2: DIC values and posterior summary of key parameters under different model assumptions for surface temperature data. The credible intervals in the last column are derived under $\sigma^{2}_{1}=\sigma^{2}_{2}$.
Figure 8 shows the posterior mean of $\mu_{1}$ and $\mu_{2}({\bf s},t)$ for the temperature series at each latitude, separated at the posterior mode of $\tau({\bf s}).$ Our method detects changepoints for all latitudes in the range $(56^{\circ}S-49^{\circ}N)$ except for latitudes $18.5^{\circ}S,26.5^{\circ}N,$ and $34.0^{\circ}N.$ No changepoints were detected in the southern $(86^{\circ}S-63.5^{\circ}S)$ and northern $(56.5^{\circ}N-86.5^{\circ}N)$ ends of the globe. The posterior probability of having no changepoint ranges from 0 to 0.0112 for detected latitudes and 0.7955 to 0.9986 for the undetected latitudes. The detected changepoints range from Sep 1991 - May 1992. See Table 3 in the Supplement for further details.
The magnitude of mean shift shown in Figure 8 is not true to scale since we normalized the data to make the variance constant for all locations in the preprocessing step. Figure 9 shows the posterior distribution of the actual mean shift at the detected latitudes, after back-transforming to match the scale of the original data. While there were more changepoints detected in the southern hemisphere, the amount of change is smaller compared to the northern hemisphere, and the change in global temperature is mostly driven by the latitudes $41.5^{\circ}N$ and $49^{\circ}N$. The average mean change in the southern hemisphere, northern hemisphere, and the globe is $-0.136^{\circ}C,-0.205^{\circ}C,$ and $-0.191^{\circ}C,$ respectively. Our findings are consistent with the existing studies (Self et al., 1996), which report that the Mt. Pinatubo eruption led to a global cooling of $0.4^{\circ}C$ between years $1991$ and $1993$, driven mostly by a cooling of $0.5^{\circ}C$ in Northern Hemisphere. The estimated change amounts returned by out model are smaller in magnitude, which is expected since our method only measures the amount of mean shift at the time of the changepoint as opposed to the average change over two years. The negative $\gamma_{1F}$ estimate also indicates that the temperature continues to cool down after the changepoint.
Fig. 8: Posterior mean of $\mu_{1}$ and $\mu_{2}({\bf s},t)$ for surface temperature, separated at the posterior mode of $\tau({\bf s})$ (red lines). The black lines are latitudinal mean temperature series after preprocessing. The posterior distribution of the changepoints are overlayed at the bottom of each plot. No changepoints were detected at $86.0^{\circ}$ $63.5^{\circ}$, $18.5^{\circ}$, $26.5^{\circ}$, $34.0^{\circ}$, and $56.5^{\circ}$ $86.5^{\circ}$.
Fig. 9: Posterior distribution of $\gamma_{0}({\bf s})$ (a) by latitude and (b) averaged over hemisphere after backtransforming to original scale. The red line on (a) marks the equator.
Our findings from the analysis of stratospheric AOD and surface temperature data are consistent with the previous studies in general. However, our work complements the previous studies by providing quantitative measures of the impact of the Pinatubo eruption on aerosol and climate, as well as specifying the spatial pattern of the impact.
6 Discussion
Motivated by need to quantify of the impact of the Mt. Pinatubo eruption on global climate and specify its spatial pattern, we proposed a Bayesian framework to simultaneously detect and estimate changepoint for spatio-temporal data. Our approach can identify which spatial locations have a changepoint and, at the same time, allows the changepoints and the magnitude of change to vary spatially, as opposed to assuming that all locations experience the same amount of change at the same time point. Furthermore, our model respects the diffusion pattern of the impact from an SAI event and takes advantage of spatial correlation in changepoint detection and estimation. In particular, our method ensures that changepoints at locations other than the event origin occur strictly after the initial changepoint. This feature helps to prevent detecting changepoints caused by unrelated events, providing a more focused analysis of our desired atmospheric event. The validity and effectiveness of our approach are demonstrated through simulations. By applying our method to AOD and surface temperature data, we successfully captured the spatial patterns of changepoints, which reveals the progression of injected aerosol and its spatially heterogeneous impact on climate.
For convenience, we have opted for the exponential covariance function, a stationary model, to address spatial correlation in the data. However, given our large spatial domain, it may be more realistic to consider a nonstationary covariance function, as suggested by strategies in Shand and Li (2017). We anticipate that computations would be more expensive when considering nonstationary spatial correlation. Additionally, we by no means imply a causal relationship between the Mt. Pinatubo eruption and the detected changepoints. These changepoints could be due to other climate and weather events, such as the El Niño effect, or other internal variations. Our aim in this article is to capture a plausible impact from the Mt. Pinatubo volcanic eruption. Separating Mt. Pinatubo’s impact from other scenarios and establishing a pure causal relationship between the volcanic eruption and climate impact is nontrivial and is a topic we reserve for future studies.
Funding Statement: Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525. SAND2024-XXXX.
7 Appendix
7.1 MCMC sampling
We write the $\boldsymbol{\mu}_{\tau}$ as
$\boldsymbol{\mu_{\tau}}$ $=\boldsymbol{\mu}_{F}+\boldsymbol{\mu}_{R}\circ\mathbf{1_{\tau}^{+}}$ (11)
$=\mathbf{Z}_{\tau}\boldsymbol{\alpha}+\boldsymbol{\mu}_{R}\circ\mathbf{1_{\tau}^{+}}$ (12)
where $\boldsymbol{\mu}_{R}=\mathbf{\gamma_{0R}}+\mathbf{\gamma_{1R}}\cdot(\mathbf{t}\otimes\mathbf{1_{N}}-\mathbf{1_{M}}\otimes\boldsymbol{\tau})$, $\mathbf{Z}_{\tau}=\begin{bmatrix}\mathbf{1_{MN}}&\mathbf{1_{\tau}^{+}}&(t\otimes\mathbf{1}_{N}-\mathbf{1}_{M}\otimes\boldsymbol{\tau})\circ\mathbf{1_{\tau}^{+}}\end{bmatrix},$ and each component of $\boldsymbol{\alpha}=\begin{bmatrix}\alpha_{0}&\gamma_{0}&\gamma_{1}\end{bmatrix}^{T}$ represents the pre-changepoint mean, amount of mean shift, and post-changepoint temporal trend, respectively.
  1. Step 1 :
    Sample $\boldsymbol{\mu}_{\tau}$ without conditioning on $\mathbf{U}.$
    1. (i)
      Given prior $\boldsymbol{\alpha}\sim N(0,s^{2}_{a}I_{3})$, the full conditional for $\boldsymbol{\alpha}$ with $\mathbf{U}$ integrated out is
      $\boldsymbol{\alpha}$ $\mid\mathbf{Y},\boldsymbol{\tau},\boldsymbol{\mu}_{R},\theta\sim N(\boldsymbol{\mu_{\alpha}},\boldsymbol{\Sigma_{\alpha}}),$
      where
      $\boldsymbol{\Sigma_{\alpha}}$ $=\left({\mathbf{Z}^{T}(\sigma^{2}_{1}\mathbf{I}_{MN}+\boldsymbol{\Sigma_{U}})^{-1}\mathbf{Z}}+\frac{\mathbf{I}_{3}}{s^{2}_{a}}\right)^{-1},$
      $\boldsymbol{\mu_{\alpha}}$ $=\boldsymbol{\Sigma_{\alpha}}\mathbf{Z}^{T}(\sigma^{2}_{1}\mathbf{I}_{MN}+\boldsymbol{\Sigma_{U}})^{-1}(\mathbf{Y}-(\boldsymbol{\epsilon_{\gamma}}+\boldsymbol{\mu}_{R})\circ\mathbf{1_{\tau}^{+}}).$
    2. (ii)
      Given prior $\tau_{0}\sim Cat(\boldsymbol{\pi}=(\pi_{1},\pi_{2},\ldots,\pi_{M})),$ the full conditional for $\tau_{0}$ with $\mathbf{U}$ integrated out is
      $\tau_{0}$ $\mid\mathbf{Y},\boldsymbol{\mu_{\tau}},\boldsymbol{\tau},\theta\sim Cat(\boldsymbol{\widetilde{\pi}}=(\widetilde{\pi}_{1},\ldots,\widetilde{\pi}_{M})),$
      where
      $\widetilde{\pi}_{k}$ $\propto\exp\left(-\frac{1}{2}(\mathbf{Z}_{\tau}\boldsymbol{\alpha}+(\boldsymbol{\epsilon_{\gamma}}+\boldsymbol{\mu}_{R})\circ\mathbf{1_{\tau}^{+}})^{T}(\sigma^{2}_{1}\mathbf{I}_{MN}+\boldsymbol{\Sigma_{U}})^{-1}(\mathbf{Z}_{\tau}\boldsymbol{\alpha}+(\boldsymbol{\epsilon_{\gamma}}+\boldsymbol{\mu}_{R})\circ\mathbf{1_{\tau}^{+}})\right).$
    3. (iii)
      See section 7.1.1 on how to sample $\boldsymbol{\Delta}$ and $\boldsymbol{\mu}_{R}$
  2. Step 2 :
    Sample $\mathbf{U}.$
    1. (i)
      The full conditional for $\mathbf{U}$ is
      $\mathbf{U}$ $\mid\cdot\sim N(\boldsymbol{\widetilde{\mu}_{U}},\boldsymbol{\widetilde{\Sigma}_{U}}),$
      where
      $\boldsymbol{\widetilde{\Sigma}_{U}}$ $=\frac{\mathbf{I}_{MN}}{\sigma^{2}_{1}}+\boldsymbol{\Sigma_{U}}^{-1},$
      $\boldsymbol{\widetilde{\mu}_{U}}$ $=\boldsymbol{\widetilde{\Sigma}_{U}}^{-1}\left(\frac{\mathbf{Y}-\mathbf{Z}_{\tau}\boldsymbol{\alpha}-(\boldsymbol{\epsilon_{\gamma}}+\boldsymbol{\mu}_{R})\circ\mathbf{1_{\tau}^{+}})}{\sigma^{2}_{1}}\right).$
  3. Step 3 :
    Sample $\theta.$
    1. (i)
      Given priors $\sigma^{2}_{1}\sim IG(a_{1},b_{1}),\sigma^{2}_{\gamma}\sim IG(a_{2},b_{2}),\sigma^{2}_{U}\sim IG(a_{3},b_{3}),\sigma^{2}_{\Delta}\sim IG(a_{4},b_{4}),$ the full conditionals are
      $\sigma^{2}_{1}\mid\cdot$ $\sim IG\left(a_{1}+\frac{MN}{2},b_{1}+\frac{(\mathbf{Y}-\mathbf{U}-\boldsymbol{\mu_{\tau}}-\boldsymbol{\epsilon_{\gamma}}\circ\mathbf{1_{\tau}^{+}}))^{T}(\mathbf{Y}-\mathbf{U}-\boldsymbol{\mu_{\tau}}-\boldsymbol{\epsilon_{\gamma}}\circ\mathbf{1_{\tau}^{+}}))}{2}\right)$
      $\sigma^{2}_{\gamma}\mid\cdot$ $\sim IG\left(a_{2}+\frac{K}{2},b_{2}+\frac{\boldsymbol{\epsilon_{\gamma}}^{T}\boldsymbol{\epsilon_{\gamma}}}{2}\right)$
      $\sigma^{2}_{U}\mid\cdot$ $\sim IG\left(a_{3}+\frac{MN}{2},b_{3}+\frac{\mathbf{U}^{T}\mathbf{R}(\phi_{U})^{-1}\otimes\mathbf{R}(\psi_{U})^{-1}\mathbf{U}}{2}\right)$
      $\sigma^{2}_{\Delta}\mid\cdot$ $\sim IG\left(a_{4}+\frac{N}{2},b_{4}+\frac{1}{2}(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})^{T}\mathbf{R}(\psi_{\Delta})^{-1}(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})\right)$
[…]
    2. (ii)
      Given $\sigma^{2}_{\gamma_{i}}\mid\gamma_{iF}\sim TruncIG(\min=0,\max=\gamma_{iF}^{2}/9,a_{5},b_{5}),\quad i\in\{0,1\},$ the full conditionals are
      $\sigma^{2}_{\gamma_{0}}\mid\cdot$ $\sim TruncIG\left(\min=0,\max=\frac{\gamma_{0F}^{2}}{9},a_{5}+\frac{N}{2},b_{5}+\frac{\boldsymbol{\gamma_{0R}}^{T}\mathbf{R}(\psi_{\gamma_{0R}})^{-1}\boldsymbol{\gamma_{0R}}}{2}\right)$
      $\sigma^{2}_{\gamma_{1}}\mid\cdot$ $\sim TruncIG\left(\min=0,\max=\frac{\gamma_{1F}^{2}}{9},a_{5}+\frac{N}{2},b_{5}+\frac{\boldsymbol{\gamma_{1R}}^{T}\mathbf{R}(\psi_{\gamma_{1R}})^{-1}\boldsymbol{\gamma_{1R}}}{2}\right)$
    3. (iii)
      Given prior $\boldsymbol{\beta}\sim N(\mathbf{0},s^{2}_{b}\mathbf{I}_{2}),$ the full conditional for $\boldsymbol{\beta}$ is
      $\boldsymbol{\beta}$ $\sim N(\boldsymbol{\mu_{\beta}},\boldsymbol{\Sigma_{\beta}}),$
      where
      $\boldsymbol{\Sigma_{\beta}}$ $=\left(\frac{\mathbf{X}^{T}\mathbf{R}(\psi_{\Delta})^{-1}\mathbf{X}}{\sigma^{2}_{\Delta}}+\frac{\mathbf{I}_{2}}{s^{2}_{b}}\right)^{-1},$
      $\boldsymbol{\mu_{\beta}}$ $=\boldsymbol{\Sigma_{\beta}}\left(\frac{\mathbf{X}^{T}\mathbf{R}(\psi_{\Delta})^{-1}\log\boldsymbol{\Delta}}{\sigma^{2}_{\Delta}}\right).$
    4. (iv)
      The full conditional for $\boldsymbol{\epsilon_{\gamma}}$ is
      $\boldsymbol{\epsilon_{\gamma}}$ $\sim N(\boldsymbol{\widetilde{\mu}_{\gamma}},\boldsymbol{\widetilde{\Sigma}_{\gamma}}),$
      where
      $\boldsymbol{\widetilde{\Sigma}_{\gamma}}$ $=diag\left(\left(\frac{\mathbf{1_{\tau}^{+}}}{\sigma^{2}_{1}}+\frac{\mathbf{1}}{\sigma^{2}_{\gamma}}\right)^{-1}\right),$
      $\boldsymbol{\widetilde{\mu}_{\gamma}}$ $=\boldsymbol{\widetilde{\Sigma}_{\gamma}}\left(\frac{Y-\mu_{\tau}-U}{\sigma^{2}_{1}}\right)\circ\mathbf{1_{\tau}^{+}}.$
    5. (v)
      Given priors $\phi_{U}\sim Unif(l_{1},u_{1}),\psi_{U}\sim Unif(l_{2},u_{2}),\psi_{\Delta}\sim Unif(l_{3},u_{3}),$ the full conditionals are
      $f(\phi_{U}\mid\cdot)$ $\propto|\mathbf{R}(\phi_{U})|^{-N/2}\exp\left(-\frac{\mathbf{U}^{T}(\mathbf{R}(\phi_{U})^{-1}\otimes\mathbf{R}(\psi_{U})^{-1})\mathbf{U}}{2\sigma^{2}_{U}}\right)\mathds{1}(l_{1}<\phi_{U}<u_{1})$
      $f(\psi_{U}\mid\cdot)$ $\propto|\mathbf{R}(\psi_{U})|^{-M/2}\exp\left(-\frac{\mathbf{U}^{T}(\mathbf{R}(\phi_{U})^{-1}\otimes\mathbf{R}(\psi_{U})^{-1})\mathbf{U}}{2\sigma^{2}_{U}}\right)\mathds{1}(l_{2}<\psi_{U}<u_{2})$
      $f(\psi_{\Delta}\mid\cdot)$ $\propto|\mathbf{R}(\psi_{\Delta})|^{-1/2}\exp\left(-\frac{(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})^{T}\mathbf{R}(\psi_{\Delta})^{-1}(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})}{2\sigma^{2}_{\Delta}}\right)\mathds{1}(l_{3}<\psi_{\Delta}<u_{3})$
      These full conditional distributions do not have a closed form. We use a Metropolis Hastings algorithm to sample these parameters. The proposal densities are
      $q(\phi_{U}^{*}\mid\phi_{U}^{k-1})$ $\sim N(\phi_{U}^{k-1},s_{1}^{2})\mathds{1}(l_{1}<\phi_{U}^{*}<u_{1})$
      $q(\psi_{U}^{*}\mid\psi_{U}^{k-1})$ $\sim N(\psi_{U}^{k-1},s_{s}^{2})\mathds{1}(l_{2}<\psi_{U}^{*}<u_{2})$
      $q(\psi_{\Delta}^{*}\mid\psi_{\Delta}^{k-1})$ $\sim N(\psi_{\Delta}^{k-1},s_{3}^{2})\mathds{1}(l_{3}<\psi_{\Delta}^{*}<u_{3})$
      The acceptance ratios are
      $r_{\phi_{U}}$ $=\left(\frac{|\mathbf{R}(\phi_{U}^{*})|}{|\mathbf{R}(\phi_{U}^{k-1}|}\right)^{-N/2}\exp\left(-\frac{\mathbf{U}^{T}((\mathbf{R}(\phi_{U}^{*})^{-1}-\mathbf{R}(\phi_{U}^{k-1})^{-1})\otimes\mathbf{R}(\psi_{U})^{-1})\mathbf{U}}{2\sigma^{2}_{U}}\right)$
      $\phantom{=}\cdot\left(\frac{\Phi(\frac{u_{1}-\phi_{U}^{k-1}}{s_{1}})-\Phi(\frac{l_{1}-\phi_{U}^{k-1}}{s_{1}})}{\Phi(\frac{u_{1}-\phi_{U}^{*}}{s_{1}})-\Phi(\frac{l_{1}-\phi_{U}^{*}}{s_{1}})}\right)$
      $r_{\psi_{U}}$ $=\left(\frac{|\mathbf{R}(\psi_{U}^{*})|}{|\mathbf{R}(\psi_{U}^{k-1}|}\right)^{-M/2}\exp\left(-\frac{\mathbf{U}^{T}(\mathbf{R}(\phi_{U})^{-1}\otimes(\mathbf{R}(\psi_{U}^{*})^{-1}-\mathbf{R}(\psi_{U}^{k-1})^{-1}))\mathbf{U}}{2\sigma^{2}_{U}}\right)$
      $\phantom{=}\cdot\left(\frac{\Phi(\frac{u_{2}-\psi_{U}^{k-1}}{s_{2}})-\Phi(\frac{l_{2}-\psi_{U}^{k-1}}{s_{2}})}{\Phi(\frac{u_{2}-\psi_{U}^{*}}{s_{2}})-\Phi(\frac{l_{2}-\psi_{U}^{*}}{s_{2}})}\right)$
      $r_{\psi_{\Delta}}$ $=\left(\frac{|\mathbf{R}(\psi_{\Delta}^{*})|}{|\mathbf{R}(\psi_{\Delta}^{k-1}|}\right)^{-1/2}\exp\left(-\frac{(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})^{T}(\mathbf{R}(\psi_{\Delta}^{*})^{-1}-\mathbf{R}(\psi_{\Delta}^{k-1})^{-1})(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})}{2\sigma^{2}_{\Delta}}\right)$
      $\phantom{=}\cdot\left(\frac{\Phi(\frac{u_{3}-\psi_{\Delta}^{k-1}}{s_{3}})-\Phi(\frac{l_{3}-\psi_{\Delta}^{k-1}}{s_{3}})}{\Phi(\frac{u_{3}-\psi_{\Delta}^{*}}{s_{3}})-\Phi(\frac{l_{3}-\psi_{\Delta}^{*}}{s_{3}})}\right)$
7.1.1 Adaptive Metropolis Hastings
To sample $\boldsymbol{\Delta},$ we draw $\log\boldsymbol{\Delta}$ and take the exponential. The log of the full conditional distribution for $\log\boldsymbol{\Delta}$ with $\mathbf{U}$ integrated out is
$\ell(\log\boldsymbol{\Delta}\mid\cdot)\propto$ $-\frac{\log\boldsymbol{\Delta}^{T}\mathbf{R}(\psi_{\Delta})^{-1}\log\boldsymbol{\Delta}-2\log\boldsymbol{\Delta}^{T}\mathbf{R}(\psi_{\Delta})^{-1}\mathbf{X}\boldsymbol{\beta}}{2\sigma^{2}_{\Delta}}$
$-\frac{1}{2}(\mathbf{Z}_{\tau}\boldsymbol{\alpha}+(\boldsymbol{\epsilon_{\gamma}}+\boldsymbol{\mu}_{R})\circ\mathbf{1_{\tau}^{+}})^{T}(\sigma^{2}_{1}\mathbf{I}_{MN}+\boldsymbol{\Sigma_{U}})^{-1}(\mathbf{Z}_{\tau}\boldsymbol{\alpha}+(\boldsymbol{\epsilon_{\gamma}}+\boldsymbol{\mu}_{R})\circ\mathbf{1_{\tau}^{+}}).$
We sample $\boldsymbol{\Delta}=(\Delta_{s_{1}},...,\Delta_{s_{N}})^{T}$ using adaptive component-wise Metropolis Hastings algorithm Roberts and Rosenthal (2009) as follows:
  1. 1.
    Set batch size $B.$ Initialize the log proposal variances $ls_{i}$ for each spatial components.
  2. 2.
    For $k=1,...,K:$
    1. (a)
      For $i=1,...,N:$
      1. i.
        Draw $\log\Delta^{*}_{s_{i}}$ from proposal density $q(\log\Delta^{*}_{s_{i}}\mid\log\Delta^{k-1}_{s_{i}})\sim N(\log\Delta_{s_{i}}^{k-1},\exp(ls_{i}))$
      2. ii.
        Calculate acceptance ratio
        $(\displaystyle r=\exp(-$$(\log\boldsymbol{\Delta}^{*}-\log\boldsymbol{\Delta}^{k-1})^{T}\mathbf{R}(\psi_{\Delta})^{-1}(\log\boldsymbol{\Delta}^{*}+\log\boldsymbol{\Delta}^{k-1}-\mathbf{X}\boldsymbol{\beta})/2\sigma^{2}_{\Delta}+$
        $(\boldsymbol{\mu_{\tau}}^{*}+\boldsymbol{\mu_{\tau}}^{k-1}+\boldsymbol{\epsilon_{\gamma}}\circ(\mathbf{1_{\tau}^{+}}^{*}+\mathbf{1_{\tau}^{+}}^{k-1})-2(\mathbf{Y}-\mathbf{U}))^{T}(\boldsymbol{\mu_{\tau}}^{*}-\boldsymbol{\mu_{\tau}}^{k-1}+\boldsymbol{\epsilon_{\gamma}}\circ(\mathbf{1_{\tau}^{+}}^{*}-\mathbf{1_{\tau}^{+}}^{k-1})/2\sigma^{2}_{1})),$
        where
        $\boldsymbol{\Delta}^{*}$ $=(\Delta_{s_{1}}^{k},...,\Delta_{s_{i-1}}^{k},\Delta^{*}_{s_{i}},\Delta_{s_{i+1}}^{k-1},...,\Delta_{s_{N}}^{k-1})^{T},$
        and
        $\mu_{\tau}^{*}(\mathbf{s},t)$ $=\alpha_{0}+[\gamma_{0}(\mathbf{s})+\gamma_{1}(\mathbf{s})\cdot(t-\tau_{0}-\Delta^{*}(\mathbf{s}))]\cdot\mathds{1}(t>\tau_{0}+\Delta^{*}(\mathbf{s})).$
      3. iii.
        Sample $u\sim Unif(0,1).$ Set $\Delta^{k}=\begin{cases}\Delta^{*}&\text{if }r>u\\ \Delta^{k-1}&o.w.\end{cases}.$ Set $a^{k}=\begin{cases}1&\text{if }r>u\\ 0&o.w.\end{cases}.$
      ljk
    2. (b)
      If $k=jB$ for integer $j\geq 1:$
      1. i.
[…]
      2. ii.
        Update $ls_{i}=\begin{cases}ls_{i}+\Delta(j),&a_{j}>0.44\\ ls_{i}-\Delta(j),&a_{j}\leq 0.44,\end{cases}$
        where $\Delta(j)=\min(0.1,1/\sqrt{j}).$
The full conditional for $\boldsymbol{\gamma_{0R}}$ is
$\mathbf{1}_{M}\otimes\boldsymbol{\gamma_{0R}}$ $\sim N(\boldsymbol{\widetilde{\mu}_{\gamma_{0}}},\boldsymbol{\widetilde{\Sigma}_{\gamma_{0}}}),$
where
$\boldsymbol{\widetilde{\Sigma}_{\gamma_{0}}}$ $=\frac{diag(\mathbf{1_{\tau}^{+}})}{\sigma^{2}_{1}}+\mathbf{I}_{N}\otimes\boldsymbol{\Sigma_{\gamma_{0}}},$
$\boldsymbol{\widetilde{\mu}_{\gamma_{0}}}$ $=\boldsymbol{\widetilde{\Sigma}_{\gamma_{0}}}^{-1}\left(\frac{diag(\mathbf{1_{\tau}^{+}})(\mathbf{Y}-\boldsymbol{\mu_{\tau}}-\boldsymbol{\epsilon_{\gamma}})}{\sigma^{2}_{1}}\right).$
Since $diag(\mathbf{1_{\tau}^{+}})$ is not separable, there is no easy way to sample from this distribution directly. Thus, we use the adaptive component-wise Metropolis Hastings to update $\boldsymbol{\gamma_{0R}}$ instead of using a Gibbs Sampler. The steps for updating $\boldsymbol{\gamma_{0R}}$ and $\boldsymbol{\gamma_{1R}}$ are same as $\boldsymbol{\Delta},$ where the acceptance ratio for $\boldsymbol{\gamma_{0R}}$ is given by
$(\displaystyle r=\exp(-$$(\boldsymbol{\gamma_{0R}}^{*}+\boldsymbol{\gamma_{0R}}^{k-1})^{T}\boldsymbol{\Sigma_{\gamma_{0}}}^{-1}(\boldsymbol{\gamma_{0R}}^{*}-\boldsymbol{\gamma_{0R}}^{k-1})/2+$
$((\mathbf{1}_{M}\otimes(\boldsymbol{\gamma_{0R}}^{*}+\boldsymbol{\gamma_{0R}}^{k-1}))\circ\mathbf{1_{\tau}^{+}}-2\mathbf{A})^{T}((\mathbf{1}_{M}\otimes(\boldsymbol{\gamma_{0R}}^{*}-\boldsymbol{\gamma_{0R}}^{k-1}))\circ\mathbf{1_{\tau}^{+}})/2\sigma^{2}_{1}),$
where
$\mathbf{A}$ $=\mathbf{Y}-\mathbf{U}-\boldsymbol{\mu}_{F}-\boldsymbol{\gamma_{1R}}\circ(\mathbf{t}\otimes\mathbf{1_{N}}-\mathbf{1_{M}}\otimes\boldsymbol{\tau}),$
and the acceptance ratio for $\boldsymbol{\gamma_{1R}}$ is given by
$(\displaystyle r=.\exp(-$$(\boldsymbol{\gamma_{1R}}^{*}+\boldsymbol{\gamma_{1R}}^{k-1})^{T}\boldsymbol{\Sigma_{\gamma_{1}}}^{-1}(\boldsymbol{\gamma_{1R}}^{*}-\boldsymbol{\gamma_{1R}}^{k-1})/2+$
$((\mathbf{1}_{M}\otimes(\boldsymbol{\gamma_{1R}}^{*}+\boldsymbol{\gamma_{1R}}^{k-1}))\circ(\mathbf{t}\otimes\mathbf{1_{N}}-\mathbf{1_{M}}\otimes\boldsymbol{\tau})\circ\mathbf{1_{\tau}^{+}}-2\mathbf{A})^{T}$
$((\mathbf{1}_{M}\otimes(\boldsymbol{\gamma_{1R}}^{*}-\boldsymbol{\gamma_{1R}}^{k-1}))\circ(\mathbf{t}\otimes\mathbf{1_{N}}-\mathbf{1_{M}}\otimes\boldsymbol{\tau})\circ\mathbf{1_{\tau}^{+}})/2\sigma^{2}_{1}),$
where
$\mathbf{A}$ $=\mathbf{Y}-\mathbf{U}-\boldsymbol{\mu}_{F}-\boldsymbol{\gamma_{0R}}.$
7.2 Table
Latitude Posterior Mode $\mathbb{P}$(no changepoint) 95% CI
$56.0^{\circ}$ Jul 1991 0 (Jun 1991, Aug 1993)
$48.5^{\circ}$ Jul 1991 0 (Jun 1991, Mar 1992)
$41.0^{\circ}$ Aug 1991 0 (Jul 1991, Feb 1992)
$33.5^{\circ}$ Sep 1991 0 (Jul 1991, Nov 1991)
$26.0^{\circ}$ Oct 1991 0.020 (Jun 1991, Aug 1994)
$18.5^{\circ}$ Dec 1996 (No CP) 0.242 (Sep 1991, Dec 1996)
$11.0^{\circ}$ Dec 1996 (No CP) 0.979 (Dec 1995, Dec 1996)
$3.5^{\circ}$ Dec 1996 (No CP) 0.960 (Jan 1996, Dec 1996)
$4.0^{\circ}$ Dec 1996 (No CP) 0.958 (Dec 1995, Dec 1996)
$11.5^{\circ}$ Sep 1991 0 (Jul 1991, Oct 1991)
$19.0^{\circ}$ Sep 1991 0 (Jun 1991, Nov 1991)
$26.5^{\circ}$ Dec 1996 (No CP) 0.976 (Dec 1996, Dec 1996)
$34.0^{\circ}$ Dec 1996 (No CP) 0.987 (Dec 1996, Dec 1996)
$41.5^{\circ}$ Jun 1992 0 (Sep 1991, Oct 1992)
$49.0^{\circ}$ Dec 1996 (No CP) 0.817 (Mar 1992, Dec 1996)
$56.5^{\circ}$ Dec 1996 (No CP) 0.818 (Jan 1992, Dec 1996)
Table 3: Summary of detected changepoints